During a 10 year period, your annual nominal income increases from 60,000 dollars to 90,000 while the inflation rate is 2% per year. Then your real income has increased by how much?
Answer #1

Where; A = Future Amount - present amante r=rak (nominal) t-time 9000oz 60000(1+ y) lo 1.5 = (tosto lera (1:5)/10 178=1041 82So, Real income will increase by = 60000 (1.02110 - Goooo x 1.23 1 $ 73800 = Real income increase